OpenStudy (anonymous):

f(x)=csc^2(sqrt(t)) differentiate the above equation

OpenStudy (anonymous):

chain rules thrice the derivative of something cubed is 3 something squared, times the derivative of something the derivative of cosecant of something is minus cosecant cotangent of something, times the derivative of something the derivative of square root of something is one over two root something

OpenStudy (anonymous):

except now that i look more carefully it is something squared whose derivative is two something

OpenStudy (anonymous):

in other words start with the square, write \[2(\text{fill in the input here})\times \frac{d}{dx}\text{rest of input}\] then \[2(\text{fill in the input here})\times -\csc(\text{input})\cot(\text{input})\times \frac{1}{2\sqrt{x}}\]